The table below shows the results for the sum when two die are rolled Die 1 Die 2 12 3 45 6 1 2 345 67 2 3 456 78 3 4 567 89 4 5 67 8 9 10 5 678 9 10 11 6 7 89 10 1112 1. What is the probability distribution? (Enter the exact values by using fractions. Use as the division bar. Do not use spaces in your answers.) 4 P(z) 10 12 P(z) 2 What is the probability of rolling two die and getting a sum of 3? P(z 3) 3. What is the probability of rolling two die and getting a sum that is less than 12? P 12) 4. What is the probability of rolling two die and getting a sum that is between 4 and 6, inclusive? P(4 5. What is the mean and standard deviation? 6) Round to the nearest hundredth, if necessary.
Answer #1

1)

from above below is distribution of x:

x 2 3 4 5 6 7
P(x)    1/36    1/18    1/12    1/9    5/36    1/6
x 8 9 10 11 12
P(x)    5/36    1/9    1/12    1/18    1/36

2))

P(X=3)=1/18

3)P(X<12) =35/36

4)

P(4<=X<=6)=1/3

5)mean =7

std deviation =2.42